OpenStudy (zenmo):

Find a set of (a) parametric equations and (b) symmetric equations for the line through the point and parallel to the specified vector or line. (For each line, write the direction numbers as integers.) Point ( -4, 1, 0), Parallel to v = (1/2)i + (4/3)j - k

OpenStudy (zenmo):

Parametric Equations of a Line in Space: \[x = x _{1}+at, y = y _{1}+bt, and z= z _{1}+ct\]

OpenStudy (michele_laino):

for part A, we have to apply this eqaution: \[\Large X = A + tv\] whic, can be rewritten by components, like below: \[\Large \left( {\begin{array}{*{20}{c}} x \\ y \\ z \end{array}} \right) = \left( {\begin{array}{*{20}{c}} { - 4} \\ 1 \\ 0 \end{array}} \right) + t\left( {\begin{array}{*{20}{c}} {1/2} \\ {4/3} \\ { - 1} \end{array}} \right)\] where t is the parameter

OpenStudy (michele_laino):

so we have: \[\Large x\left( t \right) = - 4 + \frac{t}{2}\] similarly for y(t) and z(t)

OpenStudy (zenmo):

\[< -4, 1, 0 > + t <\frac{ 1 }{ 2 }, \frac{ 4 }{ 3 }, -1>\]

OpenStudy (zenmo):

=\[<-4+\frac{ 1 }{ 2 }t, 1+\frac{ 4 }{ 3 }t, -t>\] Is that correct for parametric equations?

OpenStudy (zenmo):

@Michele_Laino

OpenStudy (zenmo):

The solution is: x = -4 + 3t, y= 1+8t, z = -6t. It doesn't make sense?

OpenStudy (michele_laino):

your parametric equations are right!

OpenStudy (michele_laino):

other parametric equations, which are equivalent to the first ones, are: \[\Large \left( {\begin{array}{*{20}{c}} x \\ y \\ z \end{array}} \right) = \left( {\begin{array}{*{20}{c}} { - 4} \\ 1 \\ 0 \end{array}} \right) + t\left( {\begin{array}{*{20}{c}} 3 \\ 8 \\ { - 6} \end{array}} \right)\]

OpenStudy (michele_laino):

so you are right!

OpenStudy (zenmo):

Yea, I figured it out, the book just formatted the answer differently by multiplying 6 to each T to get rid of the fractions.

OpenStudy (zenmo):

so its x= -4t +3, y= 1+8t, z=-6t

OpenStudy (michele_laino):

it suffice that you change your parameter, namely: \[\Large t \to 6\tau \] where \tau is the new parameter

OpenStudy (michele_laino):

\[\Large \left( {\begin{array}{*{20}{c}} x \\ y \\ z \end{array}} \right) = \left( {\begin{array}{*{20}{c}} { - 4} \\ 1 \\ 0 \end{array}} \right) + \tau \left( {\begin{array}{*{20}{c}} 3 \\ 8 \\ { - 6} \end{array}} \right)\]

OpenStudy (zenmo):

I need a slight help on another problem on converting into symmetric equations, I already did the parametric part.

OpenStudy (michele_laino):

ok! I see your new problem